OpenStudy (anonymous):

f(t)=t^2, simplify 4[f(w-1)]+10

OpenStudy (anonymous):

\[4(f(w-1)+10=4(w-1)^2+10\] then algebra

OpenStudy (anonymous):

substitute the value of t with (w-1) f(w-1)=(w-1)^2 4[(w-1)^2]+10